# On the Optimization Dynamics of RLVR: Gradient Gap and Step Size Thresholds

### v2

@@ D.1 Trajectory-Level Analysis Lower Bound (Proof of Theorem 2)
$\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)=\boldsymbol{\phi}(q,\vec{\boldsymbol{o}}\!\,)-\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}(\cdot\mid q)}\big[\boldsymbol{\phi}(q,\vec{\boldsymbol{o}}\!\,)\big]$
@@
where the last inequality is true by the lower bound on $\eta$ of Equation˜36 and the fact that $L_{\mathrm{o}}^{1/2}\leq G_{\mathrm{o}}$. Note that the normalizing constant terms arising in $\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)$ within the definitions $\boldsymbol{g}_{q}^{+}$ and $\boldsymbol{g}_{q}^{-}$ were not computed since they cancel out in the difference $\boldsymbol{g}_{q}^{+}-\boldsymbol{g}_{q}^{-}$. Therefore, the gap condition is satisfied.
D.2 Token-Level Analysis Lower Bound (Proof of Theorem˜4)
Under the token-level formulation of Equation˜1, recall that $T_{\infty}$ denotes the maximum length of a sequence $\vec{\boldsymbol{o}}\!\,\in\mathcal{O}$. We then aim to show for any $T_{\infty}$, $L_{\mathrm{p}}$, and $G_{\mathrm{p}}$ with $G_{\mathrm{p}}\geq L_{\mathrm{p}}^{1/2}$, there exists a problem instance where we have $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$ and for $\eta$ satisfying
@@ D.2 Token-Level Analysis Lower Bound (Proof of Theorem˜4)
Similar to the proof of Theorem˜2, we’ll use a token space of $\{o_{1},o_{-1},o_{-2}\}$ and we’ll use a linear feature representation for the logit $\boldsymbol{h}_{\theta}(q,\vec{\boldsymbol{o}}\!\,_{\leq t}):\,=\langle\boldsymbol{\phi}(q,\vec{\boldsymbol{o}}\!\,_{t}),\theta\rangle$ at each layer $t\in[T_{\infty}]$. Note that the feature map only depends on the last token in the subsequence $\vec{\boldsymbol{o}}\!\,_{\leq t}$ and is the same for each layer $t$. Each feature map will be
@@
$\qquad+\frac{\delta\,G}{6}\cdot\Big(\frac{T_{\infty}}{1-J_{q}(k)}\Big)\cdot\frac{1}{1+\big\{\exp(\eta\,G_{\mathrm{p}}/6)+\exp(-\eta\,G_{\mathrm{p}}/6)\big\}\cdot\exp\big\{(\delta\eta\,G_{\mathrm{p}}/6)\cdot k\big\}}\,.$
@@
Appendix E First-Order Analysis of the Log-Odds Change
Trajectory Level:
The analysis for entire generation trajectories is presented in Lemma˜1, with the full proof in Section˜E.1.
Token Level:
The corresponding result for token-level consideration is established in Lemma˜2, with its proof located in Section˜E.2.
E.1 Proof of Lemma 1
E.1.1 Overview
We first formulate Lemma˜3 below, which establishes an equivalent expression for the difference of logarithms. The proof is deferred to Section˜E.1.2.
Lemma 3.
@@ E.1.1 Overview
The difference of logarithms from Lemma˜3 is central to our analysis of the optimization scheme. As shown in equation (45), we express this term as the difference between two log moment-generating functions (MGFs). These MGFs are derived from the conditional probabilities over the correct (positive) and incorrect (negative) solution spaces, $\mathcal{O}_{q}^{+}$ and $\mathcal{O}_{q}^{-}$, respectively.
In Lemma˜4 below, we approximate the log MGFs based on the expectations of the random variables. See Section˜E.1.3 for the proof.
Lemma 4.
For any random variable $X$, we have:
We now combine Lemmas˜3 and 4 to prove Lemma˜1. In what follows, we introduce a shorthand $\boldsymbol{L}_{\theta}:\,=\log\uppi_{\theta}$. In our analysis recall that $\theta-\theta_{\rm old}=\eta_{k}\cdot\boldsymbol{w}_{k}$ with $\|\boldsymbol{w}_{k}\|_{2}\leq 1$. We denote $\eta:\,=\eta_{k}$ and $\boldsymbol{w}:\,=\boldsymbol{w}_{k}$ for short.
We apply the bound from equation (46) in Lemma˜4 to the random variable $(\boldsymbol{L}_{\theta}-\boldsymbol{L}_{\theta_{\rm old}})(q,\vec{\boldsymbol{o}}\!\,)$. By evaluating this for the distributions corresponding to the positive space, $\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta_{\rm old}}\big(\cdot\bigm|q,\mathcal{O}_{q}^{+}\big)$, and the negative space, $\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta_{\rm old}}\big(\cdot\bigm|q,\mathcal{O}_{q}^{-}\big)$, we find that
where
To establish the final bound in Lemma˜1, we need to prove two intermediate results:
  • (i)
    $T_{1}=\Delta\mu_{q}(\uppi_{\theta_{\rm old}})\cdot\eta+\mathcal{O}(\eta^{2})\qquad$ and
  • (ii)
Analyzing term $T_{1}$:
For term $T_{1}$, ˜1 on the smoothness of function $\boldsymbol{L}_{\theta}$ implies that
Moreover, the definition (11) of gap $\Delta\mu_{q}$ leads to
From inequality (48), we also have
Finalizing the proof:
Combining our previous bounds (47), (49) and (50), we get
where the last inequality follows from the condition $\eta\leq\frac{1}{2\sqrt{L_{\mathrm{o}}}}$. This establishes inequality (28) and completes the proof of Lemma˜1.
E.1.2 Proof of Lemma˜3
Therefore, we have
and find that
Similarly, for the denominator $T_{-}$, we apply equality
It follows that
Then, substituting equations (52a) and (52b) into equation (51), we obtain precisely equation (45) stated in Lemma˜3.
E.1.3 Proof of Lemma 4
$\big|\log\mathds{E}[e^{X}]-\mathds{E}[X]\big|=\log\mathds{E}[e^{Y}]\,.$
Finally, applying Hoeffding’s lemma to this resulting log moment-generating function bounds it by at most $2\,\|X\|_{\infty}^{2}$. This establishes the result in Lemma˜4.
E.2 Proof of Lemma˜2
E.2.1 Overview
$\log\Big(\frac{J_{q}}{1-J_{q}}\Big)-\log\Big(\frac{J_{q}^{\rm old}}{1-J_{q}^{\rm old}}\Big)\\ \;\geq\;\big\{\log\big(\mathds{E}^{+}[\exp(X)]\big)-\log\big(\mathds{E}^{-}[\exp(X)]\big)\big\}-L_{\mathrm{p}}T_{\infty}\cdot\eta^{2}\,$, (53)
where the random variable $X$ is defined as
@@ E.2.1 Overview
$\log\big(\mathds{E}^{+}[\exp(X)]\big)-\log\big(\mathds{E}^{-}[\exp(X)]\big)\;\geq\;\Delta\mu_{q}(\uppi_{\theta_{\rm old}})\cdot\eta-\frac{G_{\mathrm{p}}^{2}T_{\psi_{1}}}{J_{q}^{\rm old}}\cdot\eta^{2}\,$, (55)
which holds under the condition that
$\eta\;\leq\;\frac{1}{\sqrt{2\,G_{\mathrm{p}}^{2}\cdot T_{\psi_{1}}}}\,.$
Combining the bounds from (53) and (55) directly establishes the result in Lemma˜2. We now prove these two intermediate claims in Sections˜E.2.2 and E.2.3.
E.2.2 Proof of Inequality (53)
Our starting point is the smoothness property of the log-policy, as stated in ˜2. For any prompt-response pair $(q,\vec{\boldsymbol{o}}\!\,)$, the function $\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)$ is $(L_{\mathrm{p}}\cdot T_{\infty})$-smooth with respect to $\theta$. A standard result for smooth functions is that:
E.2.3 Proof of Inequality (55)
We now prove each claim in turn.
Proof of Inequality (56):
$\log(u/v)\leq\frac{1}{p}\left(\log(pu+(1-p)v)-\log(v)\right).$
Substituting (59) into (58) completes the proof of inequality (56).
Proof of Inequality (57a):
Substituting this back into our gradient identity yields:
Dividing by $J_{q}^{\rm old}$ and multiplying by $\eta$, we find
Proof of Inequality (57b):
Then $X_{t}=\sum_{t^{\prime}=1}^{t}\xi_{t^{\prime}}$ is a martingale with respect to the filtration $\{\mathcal{F}_{t}\}$, and $X=X_{|\vec{\boldsymbol{o}}\!\,|}$. By ˜2, the increments are bounded: $|\xi_{t}|\leq G_{\mathrm{p}}\cdot\eta$.

### v3

@@ F.1 Trajectory-Level Analysis Lower Bound (Proof of Theorem 2)
$\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)=\boldsymbol{\phi}(q,\vec{\boldsymbol{o}}\!\,)-\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}(\cdot\mid q)}\big[\boldsymbol{\phi}(q,\vec{\boldsymbol{o}}\!\,)\big]\,$,
@@
where the last inequality is true by the lower bound on $\eta$ of Equation 52 and the fact that $L_{\mathrm{o}}^{1/2}\leq G_{\mathrm{o}}$. Note that the normalizing constant terms arising in $\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)$ within the definitions $\boldsymbol{g}_{q}^{+}$ and $\boldsymbol{g}_{q}^{-}$ were not computed since they cancel out in the difference $\boldsymbol{g}_{q}^{+}-\boldsymbol{g}_{q}^{-}$. Therefore, the gap condition is satisfied.
F.2 Token-Level Analysis Lower Bound (Proof of Theorem 4)
Under the token-level formulation of Equation 1, recall that $T_{\infty}$ denotes the maximum length of a sequence $\vec{\boldsymbol{o}}\!\,\in\mathcal{O}$. We then aim to show for any $T_{\infty}$, $L_{\mathrm{p}}$, and $G_{\mathrm{p}}$ with $G_{\mathrm{p}}\geq L_{\mathrm{p}}^{1/2}$, there exists a problem instance where we have $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$ and for $\eta$ satisfying
@@ F.2 Token-Level Analysis Lower Bound (Proof of Theorem 4)
Similar to the proof of Theorem 2, we’ll use a token space of $\{o_{1},o_{-1},o_{-2}\}$ and we’ll use a linear feature representation for the logit $\boldsymbol{h}_{\theta}(q,\vec{\boldsymbol{o}}\!\,_{\leq t}):\,=\langle\boldsymbol{\phi}(q,\vec{\boldsymbol{o}}\!\,_{t}),\theta\rangle$ at each layer $t\in[T_{\infty}]$. Note that the feature map only depends on the last token in the subsequence $\vec{\boldsymbol{o}}\!\,_{\leq t}$ and is the same for each layer $t$. Each feature map will be
@@
$\qquad+\frac{\delta\,G_{\mathrm{p}}}{6}\cdot\Big(\frac{T_{\infty}}{1-J_{q}(k)}\Big)\cdot\frac{1}{1+\big\{\exp(\eta\,G_{\mathrm{p}}/6)+\exp(-\eta\,G_{\mathrm{p}}/6)\big\}\cdot\exp\big\{(\delta\eta\,G_{\mathrm{p}}/6)\cdot k\big\}}\,.$
@@
Appendix G Divergence of Cumulative Gap Alignment and Convergence of Rewards
Here, we show a converse of Theorem 1(b) holds, in that convergence $\lim_{K\to\infty}J_{q}(K)=1$ in fact implies divergence of the cumulative Gap Alignment $\lim_{K\to\infty}M_{q}(K)=+\infty$. We present the result for the trajectory-level formulation of Section 3 and the results for the token-level formulation of Section 4 and the batch-level formulation of Section 5 are analogous.
Furthermore, the limiting behavior of the sum of policy gradient norms $\sum_{k=0}^{K-1}\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}$ may not correspond to the convergence of rewards, i.e. the sum of gradient norms may converge while $\lim_{K\to\infty}J_{q}(K)=1$. This shows that our notion of Gap Alignment better and more sharply characterizes the improvement in rewards by adjusting for the variability scaling in Equation 9.
Theorem 6.
Consider a case where $J_{q}(0)<1$. Assume the step sizes $\eta_{k}$ satisfy
$\eta_{k}\;\leq\;\frac{[\Delta\mu_{q}(k)]_{+}}{2\,(L_{\mathrm{o}}+8\,G_{\mathrm{o}}^{2})}\land\frac{1}{2\sqrt{L_{\mathrm{o}}}}\quad\mbox{where }[\,\cdot\,]_{+}=\max(0,\cdot)\,.$ (61)
  1. (a)
    (Convergence of Reward Implies Divergence of Cumulative Gap Alignment) If $\underset{K\to\infty}{\lim}J_{q}(K)=1$, then $\underset{K\to\infty}{\lim}M_{q}(K)=+\infty$.
  2. (b)
    (Vanishing Gradient Norms under Uniform Gap) We have:
    $\sum_{k=0}^{\infty}\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}^{2}<2L_{\mathrm{o}}\sum_{k=0}^{\infty}\frac{1-J_{q}(0)}{1-J_{q}(0)+J_{q}(0)\cdot\exp\left\{\frac{M(k)}{2}\right\}}\,.$
    Furthermore, if $J_{q}(0)\in(0,1)$ and every update direction $\boldsymbol{w}_{k}$ provides a uniform gap, $\Delta\mu_{q}(k)\geq\Delta\mu_{q}>0$ for all $k\geq 0$, a fixed step size $\eta\equiv\eta_{k}$ satisfying (61) yields $\sum_{k=0}^{\infty}\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}^{2}<\infty$.
Proof.
We first show part (a). From Lemma 1, we have
$\log\Big(\frac{J_{q}(K)}{1-J_{q}(K)}\Big)-\log\Big(\frac{J_{q}(0)}{1-J_{q}(0)}\Big)\;$ $\leq\;\sum_{k=0}^{K-1}\Delta\mu_{q}(k)\,\eta_{k}+(L_{\mathrm{o}}+8\,G_{\mathrm{o}}^{2})\,\sum_{k=0}^{K-1}\eta_{k}^{2}$
$\leq\;\frac{3}{2}\sum_{k=0}^{K-1}[\Delta\mu_{q}(k)]_{+}\,\eta_{k}$
$=\;\frac{3}{2}\,M_{q}(K)\,.$
Thus, if $J_{q}(K)\to 1$, we have $M_{q}(K)\to\infty$.
Next, we show part (b). From (9), and boundedness of the policy score (Assumption 1) we have
$\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}=J_{q}(\uppi_{\theta_{k}})\cdot\{1-J_{q}(\uppi_{\theta_{k}})\}\cdot\|\boldsymbol{g}_{q}^{+}-\boldsymbol{g}_{q}^{-}\|_{2}\leq 2G_{\mathrm{o}}\cdot(1-J_{q}(\uppi_{\theta_{k}}))\,.$
At the same time, from (29) and following the calculations of the proof of Subsection C.1.2, we have that
$1-J_{q}(k)\leq\frac{1-J_{q}(0)}{1-J_{q}(0)+J_{q}(0)\cdot\exp\left\{\frac{M(k)}{2}\right\}}\,.$
Thus, $\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}=\mathcal{O}\left(\exp\left\{-K\cdot\Delta\mu_{q}\cdot\eta/2\right\}\right)$ meaning $\sum_{k=0}^{\infty}\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}^{2}<\infty$. ∎
Remark 1.
We note the mismatch in behavior between the divergence of $\lim_{K\to\infty}M_{q}(K)$ and $\sum_{k=0}^{\infty}\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}^{2}$ also occurs outside of convergence regimes. For example, in the lower bound construction of Theorem 2 (see Appendix F), one readily deduces using (9) that $\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}$ vanishes exponentially fast so that $\sum_{k=0}^{\infty}\|\nabla_{\theta}\,J_{q}(\uppi_{\theta_{k}})\|_{2}^{2}<\infty$. At the same time, since there is a uniform gap and fixed step size, $M_{q}(K)\to\infty$ while the reward vanishes $J_{q}(K)\to 0$.
Appendix H Proof of Multi-Prompt Convergence Result
H.1 Proof Sketch of Theorem 5
We first develop the following batch analogues of Lemmas 1, 2 and 3. As most of the proof details are identical to the single-prompt setting, we only give sketches of the key modifications of these lemmas in Subsection H.2.
Lemma 5 (Batch-Level Log-Odds Decomposition).
@@ H.1 Proof Sketch of Theorem 5
$\log\left(\frac{{J}(\uppi_{\theta})}{1-{J}(\uppi_{\theta})}\right)-\log\left(\frac{{J}(\uppi_{\theta_{\rm old}})}{1-{J}(\uppi_{\theta_{\rm old}})}\right)\;=\;\log\mathds{E}_{(\vec{\boldsymbol{o}}\!\,,q)\sim P^{+}_{\theta_{\rm old}}}\big[e^{\Delta\boldsymbol{L}(\vec{\boldsymbol{o}}\!\,,q)}\big]-\log\mathds{E}_{(\vec{\boldsymbol{o}}\!\,,q)\sim P^{-}_{\theta_{\rm old}}}\big[e^{\Delta\boldsymbol{L}(\vec{\boldsymbol{o}}\!\,,q)}\big]\,$, (62)
where $\Delta\boldsymbol{L}(\vec{\boldsymbol{o}}\!\,,q):\,=\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,|q)-\log\uppi_{\theta_{\rm old}}(\vec{\boldsymbol{o}}\!\,|q)$.
Lemma 6 (Batch-Level Taylor Expansion).
Suppose that Assumption 1 holds and the step size $\eta_{k}\leq\frac{1}{2\sqrt{L_{\mathrm{o}}}}$. Then
$\bigg|\log\Big(\frac{{J}(k+1)}{1-{J}(k+1)}\Big)-\log\Big(\frac{{J}(k)}{1-{J}(k)}\Big)-{\Delta\mu}(k)\,\eta_{k}\bigg|\;\leq\;(L_{\mathrm{o}}+8\,G_{\mathrm{o}}^{2})\,\eta_{k}^{2}\,.$ (63)
Lemma 7 (Batch-Level Log-Odds Lower Bound).
Suppose Assumptions 2 and 3 hold uniformly over all prompts $q$ in the support of $\mathds{P}_{Q}$. If the step size $\eta_{k}$ satisfies
$\eta_{k}\;\leq\;\frac{1}{\sqrt{2\,G_{\mathrm{p}}^{2}\cdot T_{\psi_{1}}}}\,$, (64)
then the increment of the batch log-odds satisfies
$\log\Big(\frac{{J}(k+1)}{1-{J}(k+1)}\Big)-\log\Big(\frac{{J}(k)}{1-{J}(k)}\Big)\;\geq\;{\Delta\mu}(k)\cdot\eta_{k}-\Big(L_{\mathrm{p}}\,T_{\infty}+\frac{G_{\mathrm{p}}^{2}\,T_{\psi_{1}}}{1-{J}(k)}\Big)\cdot\eta_{k}^{2}\,.$ (65)
The proof then relies on combining the “trajectory log-odds bound” (Lemma 6) and the “token-level log-odds bound” (Lemma 7), following the same case analysis as the proof of Theorem 3 in Subsection C.2.
H.2 Proofs of Auxilliary Lemmas Needed for Batch-Level Analysis
H.2.1 Proof of Lemma 5
We compute the batch success rate under the new policy:
${J}(\uppi_{\theta})$ $\;=\;\sum_{(q,\vec{\boldsymbol{o}}\!\,)\in\mathcal{O}^{+}}\mathds{P}_{Q}(q)\,\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,|q)$
$\;=\;\sum_{(q,\vec{\boldsymbol{o}}\!\,)\in\mathcal{O}^{+}}\mathds{P}_{Q}(q)\,\uppi_{\theta_{\rm old}}(\vec{\boldsymbol{o}}\!\,|q)\cdot\exp\big\{\Delta\boldsymbol{L}(\vec{\boldsymbol{o}}\!\,,q)\big\}$
$\;=\;{J}(\uppi_{\theta_{\rm old}})\cdot\mathds{E}_{(\vec{\boldsymbol{o}}\!\,,q)\sim P^{+}_{\theta_{\rm old}}}\big[e^{\Delta\boldsymbol{L}(\vec{\boldsymbol{o}}\!\,,q)}\big]\,.$
The key step is recognizing that $\mathds{P}_{Q}(q)\cdot\uppi_{\theta_{\rm old}}(\vec{\boldsymbol{o}}\!\,|q)/{J}(\uppi_{\theta_{\rm old}})=P^{+}_{\theta_{\rm old}}(q,\vec{\boldsymbol{o}}\!\,)$ for $(q,\vec{\boldsymbol{o}}\!\,)\in\mathcal{O}^{+}$. Similarly, for the negative space:
$1-{J}(\uppi_{\theta})$ $\;=\;\{1-{J}(\uppi_{\theta_{\rm old}})\}\cdot\mathds{E}_{(\vec{\boldsymbol{o}}\!\,,q)\sim P^{-}_{\theta_{\rm old}}}\big[e^{\Delta\boldsymbol{L}(\vec{\boldsymbol{o}}\!\,,q)}\big]\,.$
Taking the ratio and then the logarithm yields (62).
H.2.2 Proof Sketch of Lemma 6
The proof follows the same structure as the proof of Lemma 1 in Subsection D.1, using Lemma 5 in place of Lemma 3 from the single-prompt analysis. From Lemma 5:
$\log\left(\frac{{J}(k+1)}{1-{J}(k+1)}\right)-\log\left(\frac{{J}(k)}{1-{J}(k)}\right)\;=\;\log\mathds{E}_{P_{k}^{+}}\big[e^{\Delta\boldsymbol{L}}\big]-\log\mathds{E}_{P_{k}^{-}}\big[e^{\Delta\boldsymbol{L}}\big]\,.$
By Lemma 4 (the log-MGF bound), for any random variable $X$:
$\big|\log\mathds{E}[e^{X}]-\mathds{E}[X]\big|\;\leq\;2\|X\|_{\infty}^{2}\,.$
Applying this to both terms:
$\log\mathds{E}_{P_{k}^{+}}\big[e^{\Delta\boldsymbol{L}}\big]$ $\;=\;\mathds{E}_{P_{k}^{+}}[\Delta\boldsymbol{L}]+\mathcal{O}(\|\Delta\boldsymbol{L}\|_{\infty}^{2})\,$,
$\log\mathds{E}_{P_{k}^{-}}\big[e^{\Delta\boldsymbol{L}}\big]$ $\;=\;\mathds{E}_{P_{k}^{-}}[\Delta\boldsymbol{L}]+\mathcal{O}(\|\Delta\boldsymbol{L}\|_{\infty}^{2})\,.$
By smoothness (Assumption 1), for any $(q,\vec{\boldsymbol{o}}\!\,)$:
$\Delta\boldsymbol{L}(\vec{\boldsymbol{o}}\!\,|q)\;=\;\eta_{k}\,\langle\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,|q),\,\boldsymbol{w}_{k}\rangle+\mathcal{O}(L_{\mathrm{o}}\,\eta_{k}^{2})\,.$
Taking expectations and using the definitions of ${\boldsymbol{g}}^{+}(k)$ and ${\boldsymbol{g}}^{-}(k)$:
$\mathds{E}_{P_{k}^{+}}[\Delta\boldsymbol{L}]-\mathds{E}_{P_{k}^{-}}[\Delta\boldsymbol{L}]\;=\;\eta_{k}\,\langle{\boldsymbol{g}}^{+}(k)-{\boldsymbol{g}}^{-}(k),\,\boldsymbol{w}_{k}\rangle+\mathcal{O}(L_{\mathrm{o}}\,\eta_{k}^{2})\;=\;{\Delta\mu}(k)\,\eta_{k}+\mathcal{O}(L_{\mathrm{o}}\,\eta_{k}^{2})\,.$
For the $\|\Delta\boldsymbol{L}\|_{\infty}$ bound, by Assumption 1:
$\|\Delta\boldsymbol{L}\|_{\infty}\;\leq\;G_{\mathrm{o}}\,\eta_{k}+\mathcal{O}(L_{\mathrm{o}}\,\eta_{k}^{2})\,.$
Combining these bounds yields (63) with the same constants as the single-prompt case.
H.2.3 Batch Policy Gradient Identity
We show here the batch quantities satisfy an analogue of Equation 9.
Lemma 8 (Relationship to Batch Policy Gradient).
Specifically:
$\nabla_{\theta}\,{J}(\uppi_{\theta})\;=\;{J}(\uppi_{\theta})\cdot\{1-{J}(\uppi_{\theta})\}\cdot\big({\boldsymbol{g}}^{+}(\uppi_{\theta})-{\boldsymbol{g}}^{-}(\uppi_{\theta})\big)\,.$ (66)
Proof.
We have
$\nabla_{\theta}\,{J}(\uppi_{\theta})$ $\;=\;\nabla_{\theta}\,\sum_{(q,\vec{\boldsymbol{o}}\!\,)\in\mathcal{O}^{+}}\mathds{P}_{Q}(q)\,\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,|q)$
$\;=\;\sum_{(q,\vec{\boldsymbol{o}}\!\,)\in\mathcal{O}^{+}}\mathds{P}_{Q}(q)\,\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,|q)\,\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,|q)$
$\;=\;{J}(\uppi_{\theta})\cdot{\boldsymbol{g}}^{+}(\uppi_{\theta})\,.$
Similarly, $\nabla_{\theta}\,\{1-{J}(\uppi_{\theta})\}=\{1-{J}(\uppi_{\theta})\}\cdot{\boldsymbol{g}}^{-}(\uppi_{\theta})$, which gives $\nabla_{\theta}\,{J}(\uppi_{\theta})=-\{1-{J}(\uppi_{\theta})\}\cdot{\boldsymbol{g}}^{-}(\uppi_{\theta})$. Combining these two expressions yields (66). ∎
H.2.4 Proof Sketch of Lemma 7
The proof follows the same structure as the proof of the single-prompt Lemma 2 in Subsection D.2, substituting the conditional joint distributions $P_{k}^{+}$ and $P_{k}^{-}$.
Step 1: Reduction to MGF bounds.
From Lemma 5, we have the log-odds decomposition in terms of moment generating functions. By the smoothness of the log-policy (Assumption 2), we obtain an analogue of Equation 42 following the same steps as Subsection D.2.2:
$\log\Big(\frac{{J}(k+1)}{1-{J}(k+1)}\Big)-\log\Big(\frac{{J}(k)}{1-{J}(k)}\Big)\;\geq\;\big(\log\mathds{E}_{P_{k}^{+}}[e^{X}]-\log\mathds{E}_{P_{k}^{-}}[e^{X}]\big)-L_{\mathrm{p}}\,T_{\infty}\cdot\eta_{k}^{2}\,$,
where $X:\,=\eta_{k}\cdot\boldsymbol{w}_{k}\cdot\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,|q)$.
Step 2: Lower bound on log-MGF ratio.
Following the proof of Equation 44 in Subsection D.2.3, we establish:
$\log\mathds{E}_{P_{k}^{+}}[e^{X}]-\log\mathds{E}_{P_{k}^{-}}[e^{X}]\;\geq\;{\Delta\mu}(k)\cdot\eta_{k}-\frac{G_{\mathrm{p}}^{2}\,T_{\psi_{1}}}{1-{J}(k)}\cdot\eta_{k}^{2}\,.$ (67)
This requires three intermediate results:
(i) Relating conditional to unconditional MGFs: Using the weighted AM-GM inequality as in Equation 45, we have:
$\log\left(\frac{\mathds{E}_{P_{k}^{+}}[e^{X}]}{\mathds{E}_{P_{k}^{-}}[e^{X}]}\right)\;\geq\;-\frac{1}{1-{J}(k)}\log\mathds{E}_{P_{k}}[e^{X}]+\frac{1}{1-{J}(k)}\log\mathds{E}_{P_{k}^{+}}[e^{X}]\,$,
where $P_{k}$ denotes the unconditional joint distribution at iteration $k$.
(ii) Lower bound on positive term: By Jensen’s inequality, $\log\mathds{E}_{P_{k}^{+}}[e^{X}]\geq\mathds{E}_{P_{k}^{+}}[X]$. Using the batch gradient identity Equation 66 of Lemma 8 and the relationship ${\boldsymbol{g}}^{+}=(1-{J})\cdot({\boldsymbol{g}}^{+}-{\boldsymbol{g}}^{-})$ (derived from $\nabla_{\theta}\,{J}={J}\cdot{\boldsymbol{g}}^{+}$), we obtain:
$\frac{\mathds{E}_{P_{k}^{+}}[X]}{1-{J}(k)}\;=\;\frac{\eta_{k}\cdot\boldsymbol{w}_{k}\cdot{\boldsymbol{g}}^{+}(k)}{1-{J}(k)}\;=\;\eta_{k}\cdot\boldsymbol{w}_{k}\cdot({\boldsymbol{g}}^{+}(k)-{\boldsymbol{g}}^{-}(k))\;=\;{\Delta\mu}(k)\cdot\eta_{k}\,.$
(iii) Upper bound on unconditional log-MGF: This is the key step requiring the token-level structure. For any prompt $q$, define the martingale difference sequence $\xi_{t}:\,=\eta_{k}\cdot\boldsymbol{w}_{k}\cdot\nabla_{\theta}\,\log\uppi_{k}(o_{t}|q,\vec{\boldsymbol{o}}\!\,_{<t})$, so that $X=\sum_{t=1}^{|\vec{\boldsymbol{o}}\!\,|}\xi_{t}$.